# Diffusive Spreading Across Dynamic Mitochondrial Network Architectures

### v2

@@ Discussion
In this work, we develop a mathematical framework for mixing of diffusive material in dynamic networks and apply it to mitochondrial networks in cultured human cells. We examine two limiting regimes: one of static, well-connected physical networks and one of socially interacting homogeneous clusters. In physical networks, spreading rates increase with material diffusivity and network dimensionality. In three-dimensional social networks, material accumulates linearly in time and is limited either by the fusion rate between nearby clusters or the mobility of those clusters.
@@
Our quantitative calculations focus on a simplified dynamic system, with a single source of material held at fixed concentration. Realistic biological scenarios are of course likely to involve additional complications, including multiple sources (such as the mitochondrial nucleoids), non-trivial regulation at the source, or diffusive barriers within the mitochondria [63]. We also assumed that individual mitochondrial units are all identical in their fusion and fission behavior, whereas it is possible that some units remain isolated from the rest of the network, while others are more likely to engage in interactions. Furthermore, dynamic rearrangements beyond fusion and fission (such as tubule extension and branch sliding) may contribute to mixing within individual mitochondrial network components [76, 77, 78]. However, the minimal model presented here provides a basic building block from which more complicated scenarios can be constructed and which can serve as a null hypothesis for analyzing experimental observations.
@@
Mitochondrial networks are known to alter their architecture in response to changing environmental conditions, with excess glucose causing network fragmentation while starvation results in hyperfused networks [38]. A quantitative framework for material spreading in different structures provides insight on the implications of these transitions for many biologically critical processes, such as the distribution of harmful reactive oxygen species [47], the ability to isolate or complement deleterious DNA mutations [81, 80], and the tunneling of calcium ions, proton gradients, and ATP throughout the cell [82, 48, 49]
While we have focused on mitochondria, analogous problems of diffusive spreading arise in many systems, such as porous media in the geosciences [23], trophallactic networks in honeybees [4], and Alzheimers progression across connected regions in the brain [83]. By bringing together models of static physical networks and transiently interacting social networks, our results enable a quantitatively predictive link from network architecture to the rate of diffusive dispersion.
Acknowledgements.

### v3

@@ Discussion
In this work, we develop a mathematical framework for spreading of diffusive material in temporal networks and apply it to mitochondrial network structures in cultured human cells. We examine two limiting regimes: one of static, well-connected physical networks and one of socially interacting clusters. In static networks, spreading rates increase with material diffusivity and network dimensionality. In three-dimensional social networks, material accumulates linearly in time and is limited either by the fusion rate between nearby clusters or the mobility of those clusters. While the existence of these two extremes has been previously noted in models of epidemic spread among mobile agents [4], we now provide analytic approximations that span over a broad range of intermediate connectivities, encompassing mass-conserving spreading of material through a network of dynamic clusters.
@@
Our quantitative calculations focus on a simplified dynamic system, with a single source of material held at fixed concentration. Realistic biological scenarios are, of course, likely to involve additional complications, including multiple sources (such as the mitochondrial nucleoids), non-trivial regulation at the source, or diffusive barriers within the mitochondria [76, 75]. We also assumed that individual mitochondrial units are all identical in their fusion and fission behavior, whereas it is possible that some units remain isolated from the rest of the network, while others are more likely to engage in interactions. Dynamic rearrangements beyond fusion and fission (such as tubule extension and branch sliding) may also contribute to material transport within individual mitochondrial network components [106, 107, 108]. Furthermore, mitochondria have been observed to engage in directed motor-driven motion in both plant [35] and mammalian [100] cells, likely enhancing the rate of encounter in the fragmented regime (see SI Appendix). However, the minimal model presented here provides a basic building block from which more complicated scenarios can be constructed and which can serve as a null hypothesis for analyzing experimental observations.
This study focuses on dispersion through the mitochondrial population. A potentially fruitful avenue of future work would be exploring how mitochondrial material is dynamically distributed throughout the cellular space. Developing a mathematically tractable approach to this question would require establishing the extrinsic spatial dimensionality [93, 69] of mitochondrial clusters, as well as accounting for heterogeneous transport of the organelles [109, 59] and for the varying geometry of the intracellular space. Such studies could further illuminate how different mitochondrial architectures resolve the tension between cellular functions that require mitochondria to be broadly dispersed yet well-mixed in their complement of proteins, RNAs, and genetic components [49].
@@
Mitochondrial networks are known to alter their architecture in response to changing environmental conditions, with excess glucose causing network fragmentation and starvation resulting in hyperfused networks [42]. A quantitative framework for material spreading in different structures provides insight on the implications of these transitions for many biologically critical processes, such as the distribution of harmful reactive oxygen species [64], the ability to isolate or complement deleterious DNA mutations [111, 75], and the tunneling of calcium ions, proton gradients, and ATP throughout the cell [112, 22, 65]
While we have focused on mitochondria, analogous problems of diffusive spreading arise in many systems, such as porous media in the geosciences [27], trophallactic networks in honeybees [6], and Alzheimers progression across connected regions in the brain [113]. By bridging across models of static physical networks and transiently interacting social networks, our results enable a quantitatively predictive link from temporal network structure to the rate of diffusive dispersion.
Materials and Methods
Mathematical derivations
The exact solution and the fractal-continuum approximation for filling of highly connected stationary networks (Eq. 1, Fig. 2) were derived as shown in sections 1A,1B of the SI Appendix. The analytic approximation for the social network model (Eq. Fragmented network regime: spreading through transient interactions,3, Fig. 3) was derived as shown in section 1C of the SI Appendix.
Simulation methods
Agent-based spatially resolved simulations of dynamic networks consisting of spherocylindrical units undergoing fusion and fission were carried out as described in prior work [67]. Spreading of material within these dynamic networks was modeled using a finite volume method, with simulation details provided in section 1D of the SI Appendix. Simulation code and example parameter files are available at https://github.com/lenafabr/mitochondrialNetworks.
Cell culture and imaging methods
Cell culture, microscopy, mitochondrial network extraction, and measurements of dynamic parameters were carried out as described in section 1E of the SI Appendix. Specific plasmids used are provided in table S1 and extracted averaged parameters are given in table S2. The mitochondrial network model was modified to use a slab-like geometry for analysis of spreading in experimentally extracted network structures, as described in section 1E of the SI Appendix. Imaging data is available at https://doi.org/10.5061/dryad.cjsxksnkb.
Acknowledgements